OpenStudy (anonymous):

for my retrice.the sum of two numbers is 95, twice the larger increase by 3 times the smaller, give the sum of 232? pls. answer this. and explain this ...please

OpenStudy (anonymous):

assignment

OpenStudy (anonymous):

we guessed that!

OpenStudy (anonymous):

i do not understand..

OpenStudy (anonymous):

\[x+y=95\] \[2x+3y=232\]

OpenStudy (anonymous):

multiply first equation by -2 and add to second \[-2x-2y=-190\] \[2x+3y=232\] add to get \[y=232-190=42\]

OpenStudy (anonymous):

and since \[y=42\] and \[x+y=95\] we know \[x=53\]

OpenStudy (amistre64):

its gonna be the same answers as before I believe :)

OpenStudy (anonymous):

bigger number is 53, smaller is 42 hello sensei! before the flood?

OpenStudy (amistre64):

i remember when ... lol

OpenStudy (anonymous):

thank you..sensei